# Evaluating LLMs for Demographic-Targeted Social Bias Detection: A Comprehensive Benchmark Study

### v1

@@ Appendix A Data characteristics
Figure 1: Token count distribution
Figure 2: Social bias label co-occurrence.
@@
Text statistics. In Figure 1, we show the token count distribution for all the bias categories, including the unbiased texts. To obtain tokens, we use the RoBERTa tokenizer. We see that the token counts for each text instance are generally low, indicating our dataset mainly contains short text instances. However, there are some outliers, e.g., in GEN and SO, where we see around 400 token length texts. Upon inspection, we observed that such instances primarily come from the BBQ dataset, that have very long context texts.
Label statistics. In Figure 3, we visualize the label statistics in the final curated dataset. The visualizations show label imbalances in the data, highlighting the need for weighted loss for optimization and motivating future work to explore further fairness interventions to ensure equitable bias detection performance. The statistics show that our data contains more biased instances than unbiased ones. Furthermore, we see that most instances target a single demographic axis. However, many instances target two axes. Instances targeting more than two demographic axes are significantly fewer in our dataset. We provide more detailed label co-occurrence statistics in Figure 2. The figure shows that text instances target specific demographics more often. For instance, texts target RAC and GEN more often. Similarly, texts target DIS,AGE, and PHY relatively less often. Furthermore, we see that GEN co-occurs with many other demographic axes, e.g., SO, RAC, and DIS. Note that while RAC and REL appear together frequently, many of these instances simply target “Jewish identities.”
@@ B.1.  Prompting
B.2.  Fine-tuning
We fine-tune LLMs for sequence classification using HuggingFace’s transformers library Wolf et al. (2020), with a maximum input length of 512 tokens. For GPT-2 models, sequences are left-padded with the EOS token.
@@ Appendix C Additional evaluations
C.1.  Ablation study: in-context learning
@@
C.2.  Ablation study: Embedding model
We next examine how the choice of embedding model affects in-context learning performance for prompting, comparing BGE-M3 Chen et al. (2024b) and BCEmbedding Youdao (2023) for selecting in-context examples. The results are presented in Table 7.
@@ C.2.  Ablation study: Embedding model
C.3.  Bias detection of each bias class

### v2

@@ Appendix A Data characteristics
(a) Token count distribution
(b) Social bias label co-occurrence
Figure 1: Overview of token distribution and label co-occurrence.
@@
Text statistics. In Figure 1(a), we show the token count distribution for all the bias categories, including the unbiased texts. To obtain tokens, we use the RoBERTa tokenizer. We see that the token counts for each text instance are generally low, indicating our dataset mainly contains short text instances. However, there are some outliers, e.g., in GEN and SO, where we see around 400 token length texts. Upon inspection, we observed that such instances primarily come from the BBQ dataset, that have very long context texts.
Label statistics. In Figure 2, we visualize the label statistics in the final curated dataset. The visualizations show label imbalances in the data, highlighting the need for weighted loss for optimization and motivating future work to explore further fairness interventions to ensure equitable bias detection performance. The statistics show that our data contains more biased instances than unbiased ones. Furthermore, we see that most instances target a single demographic axis. However, many instances target two axes. Instances targeting more than two demographic axes are significantly fewer in our dataset. We provide more detailed label co-occurrence statistics in Figure 1(b). The figure shows that text instances target specific demographics more often. For instance, texts target RAC and GEN more often. Similarly, texts target DIS,AGE, and PHY relatively less often. Furthermore, we see that GEN co-occurs with many other demographic axes, e.g., SO, RAC, and DIS. Note that while RAC and REL appear together frequently, many of these instances simply target “Jewish identities.”
@@ B.1 Prompting
B.2 Fine-tuning
We fine-tune LLMs for sequence classification using HuggingFace’s transformers library (Wolf et al., 2020), with a maximum input length of 512 tokens. For GPT-2 models, sequences are left-padded with the EOS token.
@@ Appendix C Additional evaluations
C.1 Ablation study: in-context learning
@@
C.2 Ablation study: Embedding model
We next examine how the choice of embedding model affects in-context learning performance for prompting, comparing BGE-M3 (Chen et al., 2024b) and BCEmbedding (Youdao, 2023) for selecting in-context examples. The results are presented in Table 7.
@@ C.2 Ablation study: Embedding model
C.3 Bias detection of each bias class